Circular No. 8724

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 M1 (LINEAR) An apparently asteroidal object discovered in the course of the LIN-

EAR project (discovery observation tabulated below) has been found to be cometary following posting on the ‘NEO Confirmation Page’. P. Birtwhistle, Great Shefford, Berkshire, U.K., reports that CCD exposures with a 0.40-m f/6.0 Schmidt-Cassegrain reflector on June 22.0 UT show a small, concentrated coma of diameter 6′′ and mag 17.9 with a very thin, straight tail 50′′ long in p.a. 235^{o}. J. Young reports that CCD frames taken on June 22.3 UT with the Table Mountain 0.61-m f/16 Cassegrain reflector shows an 8′′ coma of mag 17.5, slightly pear-shaped along the tail axis, with a tail 40′′ long and ∼ 10′′–16′′ wide spanning p.a. 225^{o}–235^{o}.

2006 UT  α_{2000}  δ_{2000}  Mag.
June 18.32118  20^{h}00^{m}54.^{s}76  +44^{o}01′01′′.4  18.4

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2006-M26.

T = 2006 Nov. 17.002 TT  ω = 100.^{o}820
Ω = 236.028 }2000.0
q = 4.09039 AU  i = 55.944

COMETS C/1997 B5–B7, 1997 O3, 1997 V8 AND V9, 1997 X7 (SOHO)

Additional comets have been found on archival SOHO website images (cf. IAUC 8722) — all being small and stellar in appearance, of mag 7.5– 8.0, and with no tails. C/1997 O3 and C/1997 V8 were Kreutz sungrazers, C/1997 X7 was a Meyer-group object, C/1997 V9 belonged to no known group, and the rest belong to the Marsden group.

Comet 1997 UT α_{2000} δ_{2000} Inst. F MPEC
C/1997 B5 Jan. 29.497 20^{h}55.^{m}5 −16^{o}51′ C2 RK 2006-C49
C/1997 B6  29.604 20 56.1 −16 49 C2 RK 2006-C49
C/1997 B7  29.647 20 56.4 −16 49 C2 RK 2006-C49
C/1997 O3 July 31.906 8 30.0 +16 55 C3 RK 2006-J44
C/1997 V8 Nov. 6.216 14 34.7 −19 27 C3 RK 2006-J44
C/1997 V9  14.240 15 13.5 −19 32 C2 RK 2006-J44
C/1997 X7 Dec. 15.413 17 32.8 −22 14 C2 RK 2006-J45
2006 June 22  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green